OpenStudy (anonymous):

The number of endangered species in a jungle decreased from 75 to 30 in one year. What is the percent of decrease in the number of endangered species? A. 40% B. 45% C. 60% D. 150%

OpenStudy (anonymous):

\[30-75=-45\] \[-45\div30=-1.50\] -1.50x100=-150%

OpenStudy (anonymous):

so D) 150%

OpenStudy (ranga):

Percentage decrease = Actual decrease / Original Amount * 100 = (75-30)/75 * 100 = ?